OpenStudy (anonymous):

solution on textbook says Let dv = (1+2x)^-2 dx -> v = - 2*(2x+1)^-1 I know - 2*(2x+1)^-1 is the integral of (1+2x)^-2 dx but I let dv equal (1+2x)^-2 dx and I don't know where's d been(I mean d in front of dv)

OpenStudy (sirm3d):

what's the integral of dv? \[\int \mathrm dv=?\]

OpenStudy (anonymous):

There's always something in fornt of dv in my opinion

OpenStudy (sirm3d):

\[\mathrm dv = (1+2x)^{-2} \mathrm dx\\\int 1\mathrm dv = \int (1+2x)^{-2} \mathrm dx\] just integrate both sides

OpenStudy (anonymous):

∫(xe^2x)/(1+2x)^2 dx

OpenStudy (anonymous):

I know I should use intergral by parts

OpenStudy (anonymous):

and I can solve it but solution on textbook confused me alot

OpenStudy (anonymous):

it saids let u = e(x)^2 , dv = 1/(1+2x)^2 dx -> du = (x*(e)^2x + e^2x) dx

OpenStudy (anonymous):

and v = - 1 / 2(1+2x) I'd like to know why du is used not in a full sentence but in a strange one

OpenStudy (anonymous):

sorry, what would you like to integrate

OpenStudy (anonymous):

I'm working on this question ∫(xe^2x)/(1+2x)^2 dx

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I can solve this problem but I don't really understand the solutions on the textbook

OpenStudy (sirm3d):

\[\Large \int \left[x e^{2x}\right] \left[\frac{1}{(1+2x)^2}\mathrm dx\right]\] the trick in the integration by parts is to separate the integrand into two parts \(u\) that can be differentiated, and \(\mathrm dv\) that can be integrated.

OpenStudy (anonymous):

UH I know that but I don't know what can a single dv do

OpenStudy (anonymous):

if your wondering about where the dv goes in the eq. well we just dont use it but we use it in the set up process, i am not really sure as to what your trying to get to, please explin further

OpenStudy (anonymous):